OpenStudy (anonymous):

simplify the radical radical 27x4y

OpenStudy (anonymous):

thats x to the 4th power

OpenStudy (anonymous):

\[\sqrt{27x^4y}\]

OpenStudy (anonymous):

you can simplify this by noting that 27 is just 3^3 \[\sqrt{3^3x^4y}\]and after that, just divide the exponents of every number and variable inside that expression by two and bringing them out, leaving the remainder inside\[\sqrt{3^2 * 3 * x^2 * x^2 * y}\] \[3x^2\sqrt{3y}\]